At 10:00 a.m, a car leaves a house at a rate of 60 mi/h. At the same time, another car leaves the same house at a rate of 50 mi/h in the opposite direction. At what time will the cars be 330 miles apart?
Answer by checkley71(8403)

60*h+50*h=330 or 110h=330 or h=330/110 or h=3 hours these two vehicles will be 330 miles apart or at 1:00 p.m.